OpenStudy (anonymous):

solve into simplest form X(sin(theta)) - Y(sin(alpha)) / sin(theta).sin(alpha) * tan(theta).tan(alpha)/ Y(tan(alpha)) - X(tan(theta)) @Physics

OpenStudy (anonymous):

yeah its like A/B* C/D

hero (hero):

Okay

hero (hero):

\[\frac{x \sin(\theta) - y \sin(\theta)}{\sin(\theta) \sin(\alpha)} \dot\ \frac{\tan(\theta) \tan(\alpha)}{y \tan(\alpha) - xtan(\theta)}\]

OpenStudy (anonymous):

hey [ X(\sin \theta) - Y(\sin \alpha) \] you got to correct this

OpenStudy (anonymous):

sorry I mean for the first one with y it is alpha not theta

OpenStudy (anonymous):

It doesn't seem to simplify nicely but here is what an online calculator did. Check picture.